the perimeter of a rectangle is 240 inches. the length exceeds the width by 76 inches. find the length and the width
Answer by QH(50)

Width= X
Length= X+76
Perimeter= 240
P=2L+2W
240=2(x+76)+2(x)
240=4x+152
240-152=4x
88/4=x
(X)=22(Width)
(X+76)=98 (Length)
*check*
240=2(98)+2(22)
240=196+44
240=240